OpenStudy (anonymous):

If the parent function is f(x)=mx+b, what is the value of the parameter m for the line passing through the point (-2,7) and (4, 3)?

OpenStudy (anonymous):

the answer is -2/3 how do you get the answer ?

OpenStudy (mathmale):

This is just a different way of asking you for the slope. m is a parameter, yes, and that parameter is the slope of the line. b is a parameter and is the y-intercept. What's the formula for the slope of a striaght line?

OpenStudy (anonymous):

y=mx+b ?

OpenStudy (mathmale):

That, EJ, is the slope-intercept form of the equation of a straight line. Our goal here is to find the value of m, the slope. That's all. so, again, what is the formula we so often use to find the slope of a line connecting two given points (which is what you have here)?

OpenStudy (anonymous):

i dont remember

OpenStudy (mathmale):

Try an Internet search. Google "slope of a straight line."

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

so i try to find the slope with those two points the l get my answer?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{-2}}\quad ,&{\color{blue}{ 7}})\quad &({\color{red}{ 4}}\quad ,&{\color{blue}{ 3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y={\color{green}{ m}}x+b\)

OpenStudy (mathmale):

Good. If you'll insert parentheses, your result will be perfect.\[m=y2-y1/x2-x1\rightarrow m=(y _{2}-y _{1})/(x _{2}-x _{1}) \]

OpenStudy (anonymous):

Thankyou for your help!

OpenStudy (mathmale):

You have two points; you can call either one point #1 and the other one point #2. You may as well go ahead and use the naming that jdoe has suggested.

OpenStudy (mathmale):

Glad to be of help.